given that the function f(x) = 4x + 1 , find a formula for f-1(x)


1
Expert's answer
2021-06-02T12:06:01-0400

Solution:

"f(x)=y=4x+1
\\\Rightarrow y-1=4x
\\\Rightarrow \dfrac{y-1}4=x
\\\therefore f^{-1}(x)=\dfrac{x-1}4"